OpenStudy (anonymous):

1)What method(s) would you choose to solve the equation: x2 + 2x - 6 = 0 A. Square roots; there is no x-term. B. Quadratic formula, graphing; the equation cannot be factored easily since the numbers are large. <<--- C. Factoring; the equation is easily factored. D. Quadratic formula, completing the square or graphing; the coefficient of x2-term is 1, but the equation cannot be factored. 2)What are the solutions of the equation? c^2 - 10c = 0 A. 1, – squareroot10 B. 0, –10 C. 0, sqrt10 D. 0, 10 <<---

OpenStudy (anonymous):

3)Use the discriminant to determine the number of real-number solutions for the equation: 8x2 + 8x + 2 = 0 A. one solution B. two solutions <<<---- C. no solutions D. infinitely many solutions 4)What are the solutions of the equation: z2 - 12z + 36 = 0? A. -6, 6 B. 6, 6 <<<---- C. 6, -6 D. -6, -6 am i right?

jimthompson5910 (jim_thompson5910):

for #1, the equation can't be factored, so that's true but the reason it can't be factored isn't because the numbers are too big the numbers -6 and 2 aren't big at all, so that's why I'm thinking B is off a bit

jimthompson5910 (jim_thompson5910):

# 2 is correct

OpenStudy (anonymous):

1) The answer is D. It can be seen that it cannot be B since the sentence is not generally regarded as true. The values of the coefficients are not large.

jimthompson5910 (jim_thompson5910):

# 3 is incorrect # 4 is correct

OpenStudy (anonymous):

Thank you !!

OpenStudy (anonymous):

Jim, I believe the discriminant implies there are two real solutions.

OpenStudy (anonymous):

Scratch that, long day at uni. Discriminant is 0, I failed to get 4x2 =8 xD

jimthompson5910 (jim_thompson5910):

for #3 D = b^2 - 4ac D = 8^2 - 4(8)(2) ... plug in a = 8, b = 8 and c = 2 D = 64 - 32*2 D = 64 - 64 D = 0 So because the discriminant is 0, this means that there is only one solution. This solution shows up twice, but it's still considered one solution.

OpenStudy (anonymous):

#1 is wrong.

OpenStudy (anonymous):

right answer is D - Quadratic formula, completing the square or graphing; the coefficient of x2-term is 1, but the equation cannot be factored.